A federal report find that a lie detector test given to truthful persons have a probability of 0.2 of suggesting that the person is deceptive. A company asks 12 job applicants to take a lie detect test. Suppose that all 12 applicants answer truthfully.
What is the probability that exactly one is being deceptive.
What is the probability that at most one is being deceptive.
What is the mean and standard deviation of this distribution.
I am having trouble with these questions, thanks in advanced

Answer by ewatrrr(24785)

p(test suggesting person is deceptive) = .2, n = 12
1)P(x=1) = binompdf(12, .2, 1) = .2062 0r 20.62%
2)P(at most one) = 1 - P(none) = 1 - (.8)^12 = .0687 0r 6.87%
3. mean = .2*12 = 2.4, SD = sqrt(2.4*.8)